OpenStudy (anonymous):

~*~*~*~~help please!!! fan n medal!!~**~*~*~ Find the equation of the circle that has a diameter with endpoints located at (7, 3) and (7, –5).

OpenStudy (anonymous):

(1, -5) and (5, -1) the mid point is the co ordinates of the center x=(1+5)/2 =3 y=(-5-1)/2=-3 center(3,-3)

OpenStudy (anonymous):

radius =

OpenStudy (anonymous):

(3,-3)(5, -1)

OpenStudy (anonymous):

radius =

OpenStudy (anonymous):

=

OpenStudy (anonymous):

2.828?

OpenStudy (anonymous):

=

OpenStudy (anonymous):

Equation of circle=